A 2.0 µF capacitor is charged to 12 V. The voltage supply is removed and then a 4.0 µF capacitor is fitted in parallel with the 2.0 µF one. Calculate the charge stored in the 2.0 µF capacitor finally.
1
Expert's answer
2020-11-27T14:07:00-0500

Voltage across "2\mu F" capacitor ="12V"


Charge stored in it "(Q) =C\times V=2\times12=24\mu C"


When the source is removed and a "4\mu F" is connected across, the charge "Q" is conserved.


Hence "24\mu C" has to be shared between "4\mu F" and "2\mu F"


As both the capacitors are connected across, let the voltage be "V" across both.


Let charge on "2\mu F" be "Q1"

Let charge on "4\mu F" be "Q2"


We have, "Q=Q1+Q2=2V+4V=24."


"6V=24 \implies V=4V"


Charge stored in "2\mu F" finally "=2\times 4=8\mu C" .